OpenStudy (anonymous):

find the limits of lim t->0 sin(cos t)/sec t

OpenStudy (anonymous):

\[\lim_{t \rightarrow 0}\sin(\cos(t))/\sec(t)\]\[\sec(t) = 1/\cos(t)\]\[\lim_{t \rightarrow 0}\sin(\cos(t))/(1/\cos(t))=\lim_{t \rightarrow 0}\sin(\cos(t))\cos(t)\]\[\lim_{t \rightarrow 0}\sin(\cos(t))\cos(t)=\sin(\cos(0))\cos(0)=\sin(1)=0\]

OpenStudy (anonymous):

\[\sin (1)\]

OpenStudy (anonymous):

but we doing special limits

OpenStudy (anonymous):

the result is 1/2.. but i dont know how they got that

OpenStudy (anonymous):

sorry, it's sin(1) which doesn't = 0 :(, 1/2 doesn't make any sense because as t->0, cos(t) = 1

OpenStudy (anonymous):

thats what i did

OpenStudy (anonymous):

but the book its making me crazy